Seismic activity continued at a low level during the period under review, with a pattern similar to that observed over the last few days. Two long-period events were recorded;but instrumental locations for these events could not be calculated because of their small magnitude. A number of rockfall signals were also observed. Broadband tremor was recorded intermittently at the Gages station at a low amplitude. The tremor amplitude was slightly higher yesterday evening until 1800, and after 1500 today.
EDM measurements were made today at the eastern and northern triangles. The results show no changes since these triangles were last measured.
Although no visual observations of the dome were possible today, the seismicity and EDM measurements suggest that the dome continues to grow at roughly the same rate as previous days. There may have been a slight increase in venting which may be attributed to the change in weather. The situation therefore remains stable and we are expected to maintain the same level of alert (Orange I) unless there is a drastic change in the level of activity.
Ms. Joan Latchman of the SRU joined the scientific team at the observatory on March 2. She will be assisting with the daily processing and interpretation of the seismological information.